CORONA infected people in a small town grows proportional to its existing
patients. Initially the number of infected people was 50000 and the COVID-19
positive people grow at a rate of 4% per week. Formulate the suitable
mathematical model and solve the following cases by Modified Euler’s method.
a) The number of patients after 3 weeks.
b) How long it will take to double the patients.
1
Expert's answer
2020-06-11T19:09:31-0400

CORONA infected people in a small town grows proportional to its existing patients. 


"N(t)=N(0)e^{rt}"

Given "N(0)=50000, r=0.04"

a)


"N(3)=50000e^{0.04(3)}=56475 (people)"

"N(3)=56475\ people"


b)


"N(t_1)=2N(0)"

"2N(0)=N(0)e^{rt_1}"

"e^{rt_1}=2"

"rt_1=\ln 2"

"t_1={\ln 2\over r}"

"t_1={\ln 2\over 0.04}"


"N(17)=50000e^{0.04(17)}="

"=98694\ people<2\times50000\ people"

"N(18)=50000e^{0.04(18)}="

"=102722\ people>2\times50000 \ people"

"t_1=18\ weeks"